News by Date
News by Category
Solar System
Stars
White Dwarfs
Supernovas
Neutron Stars
Black Holes
Milky Way Galaxy
Normal Galaxies
Quasars
Galaxy Clusters
Cosmology/Deep Field
Miscellaneous
Press Resources
Status Reports
Press Advisories
Image Releases
Release Guidelines
Image Use Policy
Web Shortcuts
Chandra Blog
RSS Feed
Image Use Policy
Questions & Answers
Glossary of Terms
Operations CXO Status Report

Friday 21 September 2018 9.00am EDT

During the last week Chandra completed the observing schedule as planned.

A real-time procedure was executed on Sep 14 to de-activate the EPHIN and EPHIN Input/Output (EIO) units. Following the initial spontaneous reset on Aug 20, the unit subsequently underwent a number of additional resets and displayed unstable behavior, likely related to operating at temperatures well above the unit qualification limits. Given that EPHIN was no longer used operationally as a ground or on-board safing monitor, the decision was made to take no further action toward restoring functionality and instead to de-activate the units.

The schedule of targets for the next week is shown below.



------------------------------------------
        Radiation Belts                     Sep 24
        SDSSJ132457.29+4802    ACIS-S
        ClGJ104513+580833      ACIS-I
        LSBCF612-01            ACIS-S
        CHIPS0005-2758         ACIS-I
        Abell2219              ACIS-I       Sep 25
        Abell98                ACIS-I       Sep 26
        Radiation Belts                     Sep 27
        lCar                   ACIS-I
        PSZ2G113.91-37.01      ACIS-I
        Abell98                ACIS-I       Sep 28
        Abell2219              ACIS-I
        Radiation Belts                     Sep 29
        PSZ2G028.63+50.15      ACIS-I
        VHSJ2220-5618          ACIS-S
        Perseusr200NW          ACIS-I       Sep 30
        Kapteyn'sStar          ACIS-S
        Abell98                ACIS-I
        SN2013ej               ACIS-S

------------------------------------------

All spacecraft subsystems continued to support nominal operations.

Return to Status Reports